# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 3.2 Simplicial Nodes in the Associated Graph and Partial Solutions
  • Case II(c): There is a block $p_{l}$ in $S_{0}$ along its left boundary and a block $p_{r}$ in $S_{0}$ along its right that are both not covered by any rec-pack in $\mathcal{R}$ (Figure 6(b)). Let $p_{ll}$ be the block immediately to the left of $p_{l}$ and $p_{rr}$ be the block immediately to the right of $p_{r}$. The blocks $p_{ll}$ and $p_{rr}$ must be uncovered blocks lying inside $P$ (as we assumed the vertical sides of $S_{0}$ to not overlap with the vertical sides of any rec-pack in $\mathcal{R}$ or any vertical polygon edges). Clearly $S_{l}$ covers $p_{ll}$ and $S_{r}$ covers $p_{rr}$. However if $p$ is not on the left boundary of $S_{0}$, then $p$ is covered by $S_{r}$ (which also covers $p_{rr}$). This implies that $p$ and $p_{rr}$ are adjacent in $G^{\mathcal{R}}(P)$. Otherwise, if $p$ is on the left boundary of $S_{0}$, $p$ cannot be on the right boundary of $S_{0}$ (as $S_{0}$ is not a unit square). This implies that $p$ is covered by $S_{l}$ (which also covers $p_{ll}$) and this means that $p$ and $p_{ll}$ are adjacent in $G^{\mathcal{R}}(P)$. Therefore in either case there is some node $p^{\prime}$ which is not covered by $S_{0}$ but adjacent to $p$ in $G^{\mathcal{R}}(P)$. This contradicts the definition of $S_{0}$ and therefore such a case does not arise.
This completes the proof that there is a maximal square $S_{0}$ with one corner in $C_{y}$ which covers $p$ and all its neighbours in $G^{\mathcal{R}}(P)$ (hence an unambiguous square given $\mathcal{R}$), provided that horizontal sides of $S_{0}$ overlap with some horizontal polygon edges. Moreover, if vertical sides of $S_{0}$ overlap with vertical polygon edges, then a similar argument implies the existence of an unambiguous square which has a corner in $C_{x}$. This completes the proof.
We now state another result which helps us detect unambiguous squares later on.
@@
Given an orthogonal polygon $P$, let $\mathcal{R}$ be a partial solution of OPCS where $\mathcal{E}(\mathcal{R})$ does not completely cover $P$. Let $S$ be any valid square. We define $D_{x}$ and $D_{y}$ as a set of lattice points as follows.
$D_{x}$ :=$:=\{(x,y)\in\mathbb{Z}^{2}|\exists x^{\prime}\text{ which is an x-coordinate of a vertex of }P\text{ or a corner of a}$
@@
$D_{y}$ :=$:=\{(x,y)\in\mathbb{Z}^{2}|\exists y^{\prime}\text{ which is a y-coordinate of a vertex of }P\text{ or a corner of a}$
@@
Let $p$ be a node in $G^{\mathcal{R}}(P)$ where $p\in S$ and $p$ has a neighbour $p_{1}\notin S$ in $G^{\mathcal{R}}(P)$. Then there is a neighbour $p^{\prime}\notin S$ of $p$ such that $p,p^{\prime}$ are covered by a maximal square $S^{\prime}$ having a corner in $D_{x}\cup D_{y}$.
In other words, $D_{x}$ and $D_{y}$ are $x$- and $y$- coordinates appearing in $P$, $\mathcal{R}$ or $S$ as well as the coordinates which differ by at most $1$ from coordinate values in the above set and intersect with $y$ and $x$ coordinates of the polygon vertices.
Proof 3.17.
We observe that all the blocks in the axis-parallel rectangle with $p$ and $p_{1}$ lying on opposite corners lie inside $P$ and are also adjacent to $p$ in $G(P)$ (any square covering $p_{1}$ and $p$ also covers these blocks). Consider we have a token on $p_{1}$ and we move the token to the block immediately below/above it, whichever decreases the difference of the x-coordinate with $p$, if the block is not covered by any rec-pack in $\mathcal{R}\cup\{S\}$. If this is not possible or if the difference of $x$-coordinates is already 00, we move the token to the block immediately to the left/right of it, whichever decreases the difference of the y-coordinate with $p$, if the block is not covered by any rec-pack in $\mathcal{R}\cup\{S\}$. We continue to execute these moves till (i) the difference in $x$-coordinates is 00 and a move to reduce the difference in the $y$-coordinates is not possible because of a covered block, (ii) the difference in $y$-coordinates is 00 and a move to reduce the difference in the $x$-coordinates is not possible because of a covered block, or (iii) a move to reduce the difference in the $x$-coordinates is not possible because of a covered block and a move to reduce the difference in the $y$-coordinates is not possible because of a covered block. Note that the token stays in the axis-parallel rectangle with $p$ and $p_{1}$ lying on opposite corners. Let $p^{\prime}$ be the block where the token finally ends up, then $p^{\prime}$ must be a node in $G^{\mathcal{R}}(P)$; and $p^{\prime}$ must be adjacent to $p$ in $G^{\mathcal{R}}(P)$. We consider the following cases.
@@
Without loss of generality, we assume $p^{\prime}$ has a larger $x$-coordinate and a larger $y$-coordinate than $p$. Let $S^{\prime}$ be a maximal square covering both $p$ and $p^{\prime}$. Assume that (Lemma 3.6) $S^{\prime}$ has horizontal sides overlapping with horizontal polygon edges (a similar argument can be made for a vertical overlap). We keep moving $S^{\prime}$ to the left until:
  • (i) it gets obstructed by a polygon edge to its left
  • (ii) overlaps with some horizontal polygon edge at just a point
  • (iii) $p^{\prime}$ is on the right boundary of the square
Let this new square be $S^{\prime\prime}$. In the first two sub-cases, a horizontal side of $S^{\prime\prime}$ overlaps with some polygon edge (possibly at just a point). In the third case, since the process stopped with the token on $p^{\prime}$, both blocks immediately to the left and the bottom of $p^{\prime}$ are covered by some rec-pack in $\mathcal{R}\cup\{S\}$, and hence $S^{\prime\prime}$ has its right side one unit to the right of a vertical side of some edge of rec-pack in $\mathcal{R}\cup\{S\}$ (the square that covers the immediate left block of $p^{\prime}$ but not $p^{\prime}$). Therefore in all cases, $S^{\prime\prime}$ has a vertical side with x-coordinate which is either the same of that of a polygon vertex or differs by at most $1$ with x-coordinate of a corner of some rec-pack in $\mathcal{R}\cup\{S\}$. Moreover, as $S^{\prime\prime}$ has its horizontal sides overlapping with polygon edges, $S^{\prime\prime}$ has at least one corner in $D_{y}$, such that $S^{\prime\prime}$ contains both $p$ and $p^{\prime}$.
@@
Without loss of generality, let $p^{\prime}$ have a larger $x$-coordinate than $p$. Let $S^{\prime}$ be a maximal square covering both $p$ and $p^{\prime}$. If $S^{\prime}$ has its horizontal sides overlapping with polygon edges, we have an identical argument as before to prove that there is another maximal square $S^{\prime\prime}$ that covers $p$ and $p^{\prime}$ with a corner in $D_{y}$. Therefore, we only consider the case when $S^{\prime}$ has vertical sides that overlap with polygon edges (Figure 7(b)). Now, we move $S^{\prime}$ to the top until:
  • it gets obstructed by a polygon edge to its top
  • overlaps with some vertical polygon edge at just a point
@@
  • $p^{\prime}$ (and hence $p$) is on the bottom boundary of the square
Let the square obtained be $S_{t}$. In the first two cases, $S_{t}$ has a vertical edge overlapping with some vertical polygon edge (hence would have a corner in $D_{x}$, proving what we want). Therefore we only consider the third case: $p,p^{\prime}$ are on the bottom boundary of $S_{t}$. We similarly construct $S_{b}$ by moving $S^{\prime}$ to the bottom until either it gets obstructed by (i) a polygon edge to its bottom, (ii) overlaps with some vertical polygon edge at just a point, or (iii) $p^{\prime}$ (and hence $p$) is on the top boundary of the $S_{b}$. Again as the first two of these cases already achieve $S_{b}$ having a corner in $D_{x}$, we consider the third case: $p,p^{\prime}$ are on the top boundary of $S_{b}$. We observe that the bottom side of $S_{t}$ lies above or on the bottom side of $S$ (as $p,p^{\prime}$ are on the bottom boundary of $S_{t}$), and similarly the top side of $S_{b}$ lies below or on the top side of $S$. Again if $S_{t}$ has its bottom side lying on the bottom side of $S$, $S_{t}$ would have a corner in $D_{x}$ and we would be done; hence we consider that $S_{t}$ has its bottom side lying above the bottom side of $S$ (and $S_{b}$ has its top side lying below the top side of $S$). Further, the top side of $S_{t}$ must lie above or on the top side of $S$ (otherwise both the top side and the bottom edge of $S_{t}$ lie between the top side and bottom side of $S$; but $S_{t}$ covers $p^{\prime}$ which lies to the right of the right side of $S$; which means $S_{t}$ has its left side lying completely in the interior of $S$, contradicting that $S_{t}$ has vertical sides overlapping with polygon edges). However, this means if we were to vertically move a square initially positioned as $S_{t}$ (top side of $S_{t}$ above or on top side of $S$) to $S_{b}$ (top side of $S_{b}$ below top side of $S$), there would be a square $S_{1}$ with its top side lying on the top side of $S$ which contains $p$ and $p^{\prime}$. Moreover, as $S_{1}$ has its vertical sides overlapping with polygon edges (same polygon edges which were overlapping with the vertical sides of $S^{\prime}$, $S_{t}$ and $S_{b}$), $S_{1}$ has a corner in $D_{x}$.
Case II: $p$ and $p^{\prime}$ have the same $x$-coordinate.
An argument similar to the previous argument works.
@@ 3.3 Placing Rec-packs given a Partial Solution
Let $S$ be a maximal square of an orthogonal polygon $P$ with side length $d$ such that
  • The top and bottom edges of $S$ overlaps with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner of $S$ and the distance between the top right corner of $S$ and the right endpoint $e_{1}$ is more than $d$.
  • $e_{2}$ contains the bottom right corner of $S$ and the distance between the bottom right corner of $S$ and the right endpoint of $e_{2}$ is more than $d$.
  • There is a strip between $e_{1},e_{2}$, and the right side of the strip is more than $d$ distance away from the right side of $S$.
Let $S^{\prime}$ be the square generated by reflecting $S$ with respect to its right edge. Then, if there is a set of rec-packs $\mathcal{R}$ which is a partial solution of OPCS with $S\in\mathcal{R}$ and if $S^{\prime}$ does not overlap with any rec-pack in $\mathcal{R}$, then $\mathcal{R}\cup\{S^{\prime}\}$ is also a partial solution.
Proof 3.19.
Clearly $S^{\prime}$ is a valid square as it is bounded by the region in the polygon in between $e_{1}$ and $e_{2}$. Let $Y$ be strip the between $e_{1}$ and $e_{2}$.
Let $p$ be the block inside $S^{\prime}$ which is located at the top-left corner of $S^{\prime}$. Since $p$ is in the strip $Y$ defined by the region between $e_{1}$ and $e_{2}$, any valid square $S^{\prime\prime}$ covering $p$ must have side length at most $d$. Therefore, $S^{\prime\prime}$ must completely lie inside the union of the regions covered by $S$ and $S^{\prime}$.
Let $\mathcal{R}^{\prime}$ be a minimum cover of $P$ with $\mathcal{R}\subset\mathcal{R}^{\prime}$. Let $\mathcal{T}=\mathcal{R}^{\prime}\setminus\mathcal{R}$. Therefore ′′$\mathcal{R}^{\prime\prime}=\mathcal{R}\cup\mathcal{E}(\mathcal{T})$ is also a minimum covering (as (′′$\mathcal{E}(\mathcal{R}^{\prime\prime})=\mathcal{E}(\mathcal{R}^{\prime})=\mathcal{E}(\mathcal{R})\cup\mathcal{E}(\mathcal{T})$). Since $S^{\prime}$ does not overlap with any rec-pack in $\mathcal{R}$, there must be a square $S^{\prime\prime}\in\mathcal{E}(\mathcal{T})$ which covers $p$ (top-left block of $S^{\prime}$). However, $\{S,S^{\prime}\}$ covers a superset of the area that $\{S^{\prime\prime},S\}$ covers (as $S^{\prime\prime}$ has side length at most $d$). Therefore, we can remove $S^{\prime\prime}$ and add $S^{\prime}$ to $\mathcal{R}^{\prime\prime}$ (giving us ′′′=($\mathcal{R}^{\prime\prime\prime}=(\mathcal{R}^{\prime}\setminus\{S^{\prime\prime}\})\cup\{S^{\prime}\}$) and this will still be a minimum cover. As $\mathcal{R}\cup\{S^{\prime}\}\subseteq\mathcal{R}^{\prime\prime\prime}$, $\mathcal{R}\cup\{S^{\prime}\}$ is a partial solution.
Remark 3.20.
This result symmetrically holds true for all four directions.
We can easily extend this result to placing a rec-pack to the right of the given square $S$.
@@
  • The top and bottom edges of $S$ overlaps with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner of $S$ and the distance between the top right corner of $S$ and the right endpoint $e_{1}$ is more than $\eta d$.
  • $e_{2}$ contains the bottom right corner of $S$ and the distance between the bottom right corner of $S$ and the right endpoint of $e_{2}$ is more than $\eta d$.
  • There is a strip between $e_{1},e_{2}$, and the right side of the strip is more than $\eta d$ distance away from the right side of $S$.
Let $R$ be the rec-pack of width $d$ and strength $\eta$, such that the the vertical sides of $R$ are of length $d$ and the right side of the square $S$ coincides with the left side of $R$. Then, if there is a partial solution $\mathcal{R}$ of OPCS, where $S\in\mathcal{R}$ and $R$ does not overlap with any rec-pack in $\mathcal{R}$, then $\mathcal{R}\cup\{R\}$ is also a partial solution.
Proof 3.22.
We observe that $R$ lies completely inside the strip $Y$ in between $e_{1}$ and $e_{2}$. Further, we generate $\eta$ squares of the same size as $S$ as follows: $S_{1}$ is the square generated by reflecting $S$ by its right side, $S_{2}$ is the square generated by reflecting $S_{1}$ by its right side, and for any $i\in\{3,\ldots\eta\}$, $S_{i}$ is the square generated by reflecting $S_{i-1}$ by its right side. Therefore, $S_{\eta}$ is the square generated by reflecting $S_{\eta-1}$ by its right side. Then clearly, $R$ covers the exact area covered by $\{S_{1},\ldots,S_{\eta}\}$. Precisely, we have $\mathcal{E}(R)=\{S_{1},\ldots,S_{\eta}\}$. Moreover, as $R$ does not overlap with any rec-pack in $\mathcal{R}$, none of $\{S_{1},\ldots,S_{\eta}\}$ overlaps with any rec-pack in $\mathcal{R}$. This setting now allows us to use Lemma 3.18 repeatedly.
Since $S\in\mathcal{R}$, by Lemma 3.18, $\mathcal{R}\cup\{S_{1}\}$ is a partial solution. Again, as $\{S_{1}\}\in\mathcal{R}\cup\{S_{1}\}$, by Lemma 3.18, $\mathcal{R}\cup\{S_{1},S_{2}\}$ is a partial solution. Continuing, for any $i\in\{1,\ldots,\eta-1\}$, as $\mathcal{R}\cup\{S_{1},\ldots,S_{i}\}$ is a partial solution, by Lemma 3.21 $\mathcal{R}\cup\{S_{1},\ldots,S_{i},S_{i+1}\}$ is a partial solution. In particular, $\mathcal{R}\cup\{S_{1},\ldots,S_{\eta}\}=\mathcal{R}\cup\mathcal{E}(R)$ is a partial solution. This means $\mathcal{R}\cup R$ is also a partial solution.
@@
Again, this result symmetrically holds true for all four directions.
We will use Lemma 3.21 in the subsequent section to efficiently obtain a large number of squares which are also a part of a partial solution.
4 Polynomial Time Algorithm for OPCS w.r.t the Number of Polygonal Vertices
In this Section, we design an algorithm for OPCS such that the input polygon is taken in terms of its $n$ vertices. The algorithm runs in $\mathcal{O}(n^{14})$ time. We first start with some subroutines which we use for our algorithm, followed by the algorithm itself. Finally, we give the running time analysis for the algorithm.
4.1 Checking if a Set of Rec-Packs covers the Input Orthogonal Polygon
We design an algorithm that checks whether a given set of rec-packs completely covers an orthogonal polygon $P$.

### v2

@@ 4.2 Simplicial nodes in the associated Graph and partial solutions
  • Case II(c): There is a block $p_{l}$ inside $S_{0}$ along its left boundary, and a block $p_{r}$ inside $S_{0}$ along its right that are both not covered by any rec-pack in $\mathscr{R}$ (Figure 6(b)). Let $p_{ll}$ be the block immediately to the left of $p_{l}$ and $p_{rr}$ be the block immediately to the right of $p_{r}$. As the vertical sides of $S_{0}$ are assumed to not overlap with a vertical side of a rec-pack or polygon edges, the blocks $p_{ll}$ and $p_{rr}$ must be uncovered blocks lying inside $P$. $S_{l}$ covers $p_{ll}$ and $S_{r}$ covers $p_{rr}$. However if $p$ is not on the left boundary of $S_{0}$, then $p$ is covered by $S_{r}$ (which also covers $p_{rr}$). This implies that $p$ and $p_{rr}$ are adjacent in $G^{\mathscr{R}}(P)$. Similarly, if $p$ is on the left boundary of $S_{0}$, then $p$ is covered by $S_{l}$ (which also covers $p_{ll}$). Thus, $p$ and $p_{ll}$ are adjacent in $G^{\mathscr{R}}(P)$. Therefore in either case there is some node $p^{\prime}$ which is not covered by $S_{0}$ but adjacent to $p$ in $G^{\mathscr{R}}(P)$. This contradicts the definition of $S_{0}$ and therefore such a case does not arise.
Lemma 4.13 gives us a set of squares which contains at least one unambiguous square. We now require a method to detect if a given square is unambiguous or not. The following result takes us in this direction.
@@
Let $\mathscr{R}$ be a partial solution of an orthogonal polygon $P$, which does not completely cover $P$. Let $S$ be any valid square. We define $D_{x}$ and $D_{y}$ as a set of lattice points as follows.
$D_{x}$ :=$:=\{(x,y)\in\mathbb{Z}^{2}\mid\exists x^{\prime}\text{ which is an x-coordinate of a vertex of }P\text{ or a corner of a}$
@@
$D_{y}$ :=$:=\{(x,y)\in\mathbb{Z}^{2}\mid\exists y^{\prime}\text{ which is ay-coordinate of a vertex of }P\text{ or a corner of a}$
@@
Let $p$, $p_{1}$ be neighbouring nodes in $G^{\mathscr{R}}(P)$ such that $p\in S$, $p_{1}\notin S$. Then, there is a neighbour $p^{\prime}$ of $p$ in $G^{\mathscr{R}}(P)$, such that $p^{\prime}\notin S$ but $p,p^{\prime}\in S^{\star}$ for some maximal square $S^{\star}$ having a corner in $D_{x}\cup D_{y}$.
Before we start formally proving Lemma 4.17, we provide some intuition on the statement, and briefly explain why this is needed. Note that, $D_{x}$ is a collection of x-coordinates of vertices/corners in $P$, $\mathscr{R}$ or $S$ as well as the x-coordinates differing by at most $1$, intersecting with y-coordinates of the vertices of $P$. $D_{y}$ is defined symmetrically. Let $\mathcal{D}$ be the set of all maximal squares with a corner in $D_{x}\cup D_{y}$. Now, what Lemma 4.17 mentions is: if there is a neighbour of $p\in S$ outside $S$, then one of the squares in $\mathcal{D}^{\prime}$ will cover $p$ as well as something outside $\mathscr{R}\cup\{S\}$.
This will be particularly useful in detecting unambiguous squares as follows. Let $S$ be the square that is to be tested if it is an unambiguous square given $\mathscr{R}$. If $S$ contains a simplicial node $p$ such that $S$ covers all neighbours of $p$ in $G^{\mathscr{R}}(P)$, then by definition, $S$ is an unambiguous square. No square in $\mathcal{D}^{\prime}$ can cover $p$ as well as any uncovered block $q$ outside $S$, because $q$ must be a neighbour of $p$ in $G^{\mathscr{R}}(P)$, $S$ already covers all neighbours of $p$. On the other hand, if $S$ is not an unambiguous square, all blocks inside $S$ would either be covered by rec-packs in $\mathscr{R}$ or by a square in $\mathcal{D}^{\prime}$ that covers something else that is not covered by $\mathscr{R}\cup\{S\}$. We now prove Lemma 4.17.
Proof 4.18 (Proof of Lemma 4.17).
Without loss of generality, assume that the x-coordinate and the y-coordinate of $p_{1}$ are not less than the x-coordinate and the y-coordinate of $p$ respectively. Let $L$ be the axis-parallel rectangular region with $p$ and $p_{1}$ being diagonally opposite corner blocks. Any valid square containing $p$ and $p_{1}$ contains the entire region of $L$. So, all blocks lying in $L$ are neighbours of $p$ in $G^{\mathscr{R}}(P)$.
Consider placing a token on $p_{1}$. We move the token as follows:
  • If the block immediately below the token is not covered by $\mathscr{R}\cup\{S\}$, and does not have a y-coordinate less than that of $p$, then move the token one block down.
  • If the previous step is not possible, and if the block immediately to the left of the token is not covered by $\mathscr{R}\cup\{S\}$, and does not have an x-coordinate less than that of $p$, then move the token one block to the left.
  • Stop when none of these are applicable. Let the final block of the token be denoted as $p^{\prime}$.
Note that the token stays inside $L$, and therefore, $p^{\prime}$ must be a node in $G^{\mathscr{R}}(P)$. Further, $p^{\prime}$ must be adjacent to $p$ in $G^{\mathscr{R}}(P)$. Since the process terminated, the block immediately to the left of $p^{\prime}$ and the block immediately below $p^{\prime}$ are either covered with some rec-pack in $\mathscr{R}\cup\{S\}$, or lie outside the polygon $P$. We consider the following cases.
@@
Since the process stopped with the token on $p^{\prime}$, the block immediately to the left and the block immediately below $p^{\prime}$ are covered by rec-packs in $\mathscr{R}\cup\{S\}$. Let $S^{\prime}$ be a maximal square covering both $p$ and $p^{\prime}$. From Lemma 4.5, $S^{\prime}$ has either its horizontal sides or its vertical sides overlapping with polygon edges. Assume that its horizontal sides overlapping with horizontal polygon edges $e_{1}$, $e_{2}$ (a similar argument can be made for vertical sides). We keep translating $S^{\prime}$ to the left until:
  • (i) it gets obstructed by a polygon edge to its left and translating it any further would take a part of the square outside the polygon $P$
  • (ii) it touches $e_{1}$ at just a point, and translating it any further would make the square lose contact with $e_{1}$
  • (iii) it touches $e_{2}$ at just a point, and translating it any further would make the square lose contact with $e_{2}$
  • (iv) $p^{\prime}$ is on the right boundary of the square, and translating it any further would take $p^{\prime}$ outside the square
Let this translated square be $S^{\prime\prime}$. Due the the fourth condition, $S^{\prime\prime}$ contains $p$ and $p^{\prime}$. Note that since $S^{\prime\prime}$ was obtained by translating $S^{\prime}$ to the left, the y-coordinates of the corners of $S^{\prime\prime}$ will share the y-coordinates of $e_{1}$ or $e_{2}$. Therefore, the corners of $S^{\prime\prime}$ share y-coordinates of some vertices of $P$.
If the translation of $S^{\prime}$ to $S^{\prime\prime}$ terminated due to the first three conditions, then a vertical side of $S^{\prime\prime}$ must overlap with some polygon edge (possibly at just a point). Therefore, one of the corners of $S^{\prime\prime}$ shares its x-coordinate with a vertex in $P$, and also its y-coordinate with a vertex in $P$. So, $S^{\star}=S^{\prime\prime}$ satisfies the conditions of the lemma, as these coordinates appear in $D_{x}$, as well as $D_{y}$.
We now consider the fourth condition: the translation was stopped as $p^{\prime}$ is on the right boundary. Recall that both blocks immediately to the left and below $p^{\prime}$ are covered by some rec-packs $\mathscr{R}\cup\{S\}$. Therefore, $S^{\prime\prime}$ has its right side one unit to the right of a vertical side of some edge of rec-pack in $\mathscr{R}\cup\{S\}$ (the rec-pack that covers the immediate left block of $p^{\prime}$ but not $p$). Then, the x-coordinate of at least one corner of $S^{\prime\prime}$ differs by 1 from the x-coordinate of a corner of some rec-pack. Therefore, $S^{\star}=S^{\prime\prime}$ has a corner in $D_{y}$.
@@
Let $S^{\prime}$ be a maximal square covering both $p$ and $p^{\prime}$. If $S^{\prime}$ has its horizontal sides overlapping with polygon edges, we have an identical argument as before: translate $S^{\prime}$ as long as the four conditions hold true, and then the exact same steps prove that the translated square has a corner in $D_{y}$. Therefore, we only consider the case when $S^{\prime}$ has vertical sides that overlap with polygon edges $e_{1}$, $e_{2}$ (Figure 7(b)). Now, we translate $S^{\prime}$ to the top until:
  • (i) it gets obstructed by a polygon edge to its top and translating it any further would take a part of the square outside the polygon $P$
  • (ii) it touches $e_{1}$ at just a point, and translating it any further would make the square lose contact with $e_{1}$
  • (iii) it touches $e_{2}$ at just a point, and translating it any further would make the square lose contact with $e_{2}$
@@
  • (iv) $p^{\prime}$ (and hence $p$) is on the bottom boundary of the square, and translating it any further would take $p^{\prime}$ (and hence $p$) outside the square
Let the square obtained be $S_{t}$. Similar to the previous case, if the translation terminated because of any of the first three conditions, then $S_{t}$ must have a horizontal side overlapping with some horizontal polygon edge (hence would have a corner in $D_{x}$, proving what we want).
Now, consider the fourth case: $p,p^{\prime}$ are on the bottom boundary of $S_{t}$. We similarly construct $S_{b}$ by moving $S^{\prime}$ to the bottom until either it gets obstructed by (i) a polygon edge to its bottom, (ii) overlaps with $e_{1}$ or $e_{2}$ at just a point, or (iii) $p^{\prime}$ (and hence $p$) is on the top boundary of the $S_{b}$. Again as the first two of these cases already achieve $S_{b}$ having a corner in $D_{x}$, we consider the third case: $p,p^{\prime}$ are on the top boundary of $S_{b}$.
We observe that the bottom side of $S_{t}$ does not lie below the bottom side of $S$, as $p,p^{\prime}$ are on the bottom boundary of $S_{t}$. Similarly the top side of $S_{b}$ does not lie above or on the top side of $S$. Again if $S_{t}$ has its bottom side lying on the bottom side of $S$, $S_{t}$ would have a corner in $D_{x}$ and we would be done. So, we consider the case when $S_{t}$ has its bottom side lying above the bottom side of $S$ and $S_{b}$ has its top side lying below the top side of $S$.
Claim 3.
The top side of $S_{t}$ must lie above or on the top side of $S$
Proof 4.19 (Proof of claim).
Assume the contrary. Then, both the top side and the bottom edge of $S_{t}$ lie between the top side and bottom side of $S$. But $S_{t}$ covers $p^{\prime}$ which lies to the right of the right side of $S$. This means $S_{t}$ has its left side lying completely in the strict interior of $S$, contradicting that $S_{t}$ has vertical sides overlapping with polygon edges $e_{1}$, $e_{2}$.
This means, if we were to vertically move a square initially positioned as $S_{t}$ (top side of $S_{t}$ above the top side of $S$) to $S_{b}$ (top side of $S_{b}$ below the top side of $S$), there would be a square $S_{1}$ with its top side lying on the top side of $S$ which contains $p$ and $p^{\prime}$. Moreover, as $S_{1}$ has its vertical sides overlapping with polygon edges $e_{1}$, $e_{2}$, $S^{\star}=S_{1}$ has a corner in $D_{x}$.
Case II I: $p$ and $p^{\prime}$ have the same $x$-coordinate.
An argument symmetric to the previous argument follows for this case.
@@ 4.3 Placing rec-packs given a partial solution
Let $S$ be a maximal square of an orthogonal polygon $P$ with side length $d$, such that
  • the top and bottom sides of $S$ overlap with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner $a$ of $S$
  • $e_{2}$ contains the bottom right corner $b$ of $S$
  • there is a strip $Y$ between $e_{1},e_{2}$, and the right side of the strip is more than $d$ distance away from the right side of $S$.
Let $S^{\prime}$ be the square generated by reflecting $S$ with respect to its right side. Then, for any partial solution $\mathscr{R}$ containing $S$ and with no overlap with $S^{\prime}$, $\mathscr{R}\cup\{S^{\prime}\}$ is also a partial solution.
Proof 4.21.
Figure 8: Proof of Lemma 4.20
$S^{\prime}$ is a valid square as it lies completely inside $Y$. Let $p$ be the block inside $S^{\prime}$ which is located at the top-left corner of $S^{\prime}$. Consider a partial solution $\mathscr{R}$ containing $S$ but not $S^{\prime}$.
Any block $p^{\prime}$ other than $p$ in $G^{\mathscr{R}}(P)$ is a neighbour of $p$, if and only if $p^{\prime}$ is covered by $S^{\prime}$. Hence all neighbours of $p$ in $G^{\mathscr{R}}(P)$ are covered by $S^{\prime}$ and hence they must induce a clique in $G^{\mathscr{R}}(P)$. Therefore $p$ is a simplicial node in $G^{\mathscr{R}}(P)$, making $S^{\prime}$ an unambiguous square given $\mathscr{R}$. Hence $\mathscr{R}\cup\{S^{\prime}\}$ is a partial solution (Lemma 4.11).
Note that in Lemma 4.20, $S^{\prime}$ forms a rec-pack of width $d$ and strength $1$. However, if the strip $Y$ is long enough, we can repeat the same procedure and reflect $S^{\prime}$ with respect to its ride side to get $S^{\prime\prime}$. Now $S^{\prime}\cup S^{\prime\prime}$ gives us a rec-pack of strength $2$. We can keep repeating this and get rec-packs of much larger strengths, as long as the strip $Y$ permits. We formalize this in the next result.
@@
  • the top and bottom sides of $S$ overlaps with horizontal polygon edges $e_{1}$ and $e_{2}$ respectively.
  • $e_{1}$ contains the top right corner of $S$
  • $e_{2}$ contains the bottom right corner of $S$
  • there is a strip $Y$ between $e_{1},e_{2}$, and the right side of the strip is more than $\eta d$ distance away from the right side of $S$.
Let $R$ be the rec-pack of width $d$ and strength $\eta$, such that
  • the the vertical sides of $R$ are of length $d$, and the horizontal sides are of length $\eta d$
  • the right side of the square $S$ coincides with the left side of $R$
If there is a partial solution $\mathscr{R}$ containing $S$ such that $R$ does not overlap with any rec-pack in $\mathscr{R}$, then $\mathscr{R}\cup\{R\}$ is also a partial solution.
Proof 4.23.
The proof is just a repeated application of Lemma 4.20. Let $S_{1}$ be the square obtained by reflecting $S$ along its right side, let $S_{2}$ be the square obtained by reflecting $S_{1}$ along its right side, and so on, till we define $S_{\eta}$. Note that $R$ is same as the region defined by $S_{1}\cup S_{2}\cup\cdots\cup S_{\eta}$.
Applying Lemma 4.20 to $\mathscr{R}$ and $S$, we get that $\mathscr{R}\cup\{S_{1}\}$ is a partial solution. Applying it again to $\mathscr{R}\cup\{S_{1}\}$ and $S_{1}$, we get ($\mathscr{R}\cup(\{S_{1}\}\cup\{S_{2}\})$ is a partial solution. Continuing in this fashion gives us that ($\mathscr{R}\cup(\{S_{1}\}\cup\{S_{2}\}\cup\cdots\{S_{\eta}\})$. This directly implies that $\mathscr{R}\cup\{R\}$ is a partial solution.
@@
Lemma 4.20 and Lemma 4.22 holds true symmetrically for all four directions.
Notice that in Lemma 4.22, $\eta$ can be arbitrarily large. However, detecting if Lemma 4.22 is applicable does not become any more difficult for large values of $\eta$. This will be crucially used to add multiple squares as a single rec-pack to the partial solution, in a single iteration. In later sections we will show how this makes our algorithm run in polynomial time with respect to $n$, the guarantee being independent of the output.
5 Polynomial-Time Algorithm with respect to the Number of Polygonal Vertices
In this Section, we design polynomial-time algorithm for OPCS with respect to the number of vertices $n$, of the orthogonal polygon. The algorithm runs in $\mathcal{O}(n^{10})$ time. We first discuss some subroutines the algorithm algorithm uses, followed by the algorithm itself with its analysis.
5.1 Checking if a set of rec-packs covers an orthogonal polygon
The first subtask we discuss is to check whether a given set of rec-packs $\mathscr{R}$ completely covers an orthogonal polygon $P$. However, this is equivalent to checking if the area of the rec-packs in $\mathscr{R}$ is equal to the area of $P$.
However, this is exactly Klee’s measure problem in two dimensions [14], asking the area of the union of rectangles. Using Bentley’s Algorithm [7], this can be solved in time $\mathcal{O}(|\mathscr{R}|\log|\mathscr{R}|)$.